The sum of the first five terms of an arithmetic series is 25. The sum of the first seven terms is 49. What are the values of t1, t12 and S12?
Answer by richwmiller(17219)

S = ½(2t1 + (n-1)d)n
25 = ½(2t1 + 4d)5,
49 = ½(2t1+ 6d)7
t1 = 1, d = 2
S12 = ½(2 + (11)2)12=144
tn = t1 + (n - 1)d
t12= 1 + (11)2=23